OpenStudy (anonymous):

Factor Completely: 8x^3-27=0

OpenStudy (aum):

\[\large 8x^3 - 27 = (2x)^3 - (3)^3 \\ \large \text{Use: }a^3 - b^3 = (a-b)(a^2+ab+b^2) \]

OpenStudy (aum):

Are you sure the "= 0" part is in the problem? Normally they ask you to factor completely an expression. With an equation they ask you to solve for x.

OpenStudy (anonymous):

Yes the Zero is in the problem